OpenStudy (anonymous):

evaluate 9x^2y^-2 for x=3 and y=2

OpenStudy (zale101):

9*(3)^2*(2)^2 = ?

OpenStudy (zale101):

=324 Did that made sense to you?

OpenStudy (anonymous):

the last 2 is to the -2 power not positive and yeah that made since i kinda did it like that the first time and my teacher wrote to use PEMDAS

OpenStudy (zale101):

Oh sorry did see that..

OpenStudy (zale101):

9*(3)^2*(2)^-2 = 20.25

OpenStudy (anonymous):

oh and it should say for x=-3 and y=2 my bad

OpenStudy (zale101):

9(-3)^2*(2)^2 = 324

OpenStudy (anonymous):

when i did it earlier i did this 9(-3)^2 * 2^-2 = 27^2 *1/4=729* 1/4 =183.25 can you tell me where i went wrong?

OpenStudy (zale101):

First- Substitute the given x and y for the expression. Second- 9*((-3)^2)*(2^2) You first solve the parenthesis, what you did is you multiplied 9 by -3 and then you squared root it.